OpenStudy (anonymous):

Rahim sets out to cross a forest. On the first day, he completes 1/10th of the journey. On the second day, he covers 2/3rd of the distance travelled the first day. He continues in this manner, alternating the days in which he travels 1/10th of the distance still to be covered, with days on which he travels 2/3 of the total distance already covered. At the end of seventh day, he finds that 22 ½ km more will see the end of his journey. How wide is the forest?

ganeshie8 (ganeshie8):

looks hard >.<

OpenStudy (anonymous):

1/10->first day 2/3*1/10->second day so1/10+1/15=1/6

OpenStudy (anonymous):

1/6-1=5/6 remaining

OpenStudy (anonymous):

5/6*1/10=1/12,1/12-5/6=3/4-->3rd day

OpenStudy (anonymous):

3/4*2/3=1/2,1/2-3/4=1/4 -->4th day

OpenStudy (anonymous):

1/4*1/10=1/40,1/4-1/40=9/40-->5th day

OpenStudy (anonymous):

9/40*2/3=3/20,3/20-9/40=3/40-->6th day

OpenStudy (anonymous):

3/40*1/10=3/400,3/400-3/40=27/400-->7th day

OpenStudy (anonymous):

something went wrong?

OpenStudy (anonymous):

remaining should be calculated carefully.always make mistake

OpenStudy (anonymous):

@ganeshie8 u know very well to calculate remaining

OpenStudy (anonymous):

if remaining is correct then we get the answer i think

OpenStudy (anonymous):

i think 4th day is 2/3*1/12=1/18,1/18-3/4=2/3

ganeshie8 (ganeshie8):

im getting 120km

ganeshie8 (ganeshie8):

total width of forest = 120km

OpenStudy (anonymous):

yes u are right!:)

OpenStudy (anonymous):

huh....

OpenStudy (anonymous):

u take a long time to reply

ganeshie8 (ganeshie8):

say the forest is \(\large x\) km : \[ \begin{array}{c|c|c} \text{day}&\text{distance covered}& \text{total distance covered} & \text{distance remaining} \\ \hline \\ 1&\color{red}{\dfrac{1}{10}(x)}&\dfrac{1}{10}(x) &\dfrac{9}{10}(x) \\ \hline \\ 2&\color{red}{\dfrac{2}{3} . \dfrac{1}{10}(x)}& \dfrac{1}{6}(x) &\dfrac{5}{6}(x) \\ \hline \\ \end{array} \]

ganeshie8 (ganeshie8):

lets fill that table ^^

OpenStudy (anonymous):

o-o its not loaded fully

ganeshie8 (ganeshie8):

Oh must be a poblem with latex, il take a screenshot and attach

ganeshie8 (ganeshie8):

http://prntscr.com/4dcv5k

OpenStudy (anonymous):

yes i am done that upto 2nd day earlier

OpenStudy (anonymous):

3rd day balance=59/60 ahhh

OpenStudy (anonymous):

13/15 ahh

ganeshie8 (ganeshie8):

say the forest is \(\large x\) km : \[ \begin{array}{c} \text{day}&\text{distance covered}& \text{total distance covered} & \text{distance remaining} \\ \hline \\ 1&\color{red}{\dfrac{1}{10}(x)}&\dfrac{1}{10}(x) &\dfrac{9}{10}(x) \\ \hline \\ 2&\color{red}{\dfrac{2}{3} . \dfrac{1}{10}(x)}& \dfrac{1}{6}(x) &\dfrac{5}{6}(x) \\ \hline \\ 3&\color{red}{\dfrac{1}{10} . \dfrac{5}{6}(x)}& \dfrac{3}{12}(x) &\dfrac{9}{12}(x) \\ \hline \\ 4&\color{red}{\dfrac{2}{3} . \dfrac{3}{12}(x)}& \dfrac{5}{12}(x) &\dfrac{7}{12}(x) \\ \hline \\ \end{array} \]

ganeshie8 (ganeshie8):

first four days ^^

ganeshie8 (ganeshie8):

here is the screenshot http://prntscr.com/4dcy9j

OpenStudy (anonymous):

same i have done ya instead of 9/12 i made a mistake of 3/4 .then it will be correct .see at the top

OpenStudy (anonymous):

5/6*1/10=1/12,1/12-5/6=3/4-->3rd day

ganeshie8 (ganeshie8):

9/12 is same as 3/4 :)

OpenStudy (anonymous):

how u got 5/12 in total distance column

OpenStudy (anonymous):

3/4*2/3=1/2,1/2-3/4=1/4 -->4th day

OpenStudy (anonymous):

oh i went wrong

OpenStudy (anonymous):

oh i multiplied the balance into 2/3 instead 3/12 of 3rd day

OpenStudy (anonymous):

@ganeshie8 can you please help me on my questions i tagged you

OpenStudy (anonymous):

5 1/10*7/12 63/120 57/120

OpenStudy (anonymous):

is it right? for 5th day

OpenStudy (anonymous):

fast @ganeshie8 .always i make mistake in calculation........

OpenStudy (anonymous):

15/80x=45/2

ganeshie8 (ganeshie8):

\[ \begin{array}{c} \text{day}&\text{distance covered}& \text{total distance covered} & \text{distance remaining} \\ \hline \\ 1&\color{red}{\dfrac{1}{10}(x)}&\dfrac{1}{10}(x) &\dfrac{9}{10}(x) \\ \hline \\ 2&\color{red}{\dfrac{2}{3} . \dfrac{1}{10}(x)}& \dfrac{1}{6}(x) &\dfrac{5}{6}(x) \\ \hline \\ 3&\color{red}{\dfrac{1}{10} . \dfrac{5}{6}(x)}& \boxed{\dfrac{3}{12}(x)} &\dfrac{9}{12}(x) \\ \hline \\ 4& \boxed{\color{red}{\dfrac{2}{3} . \dfrac{3}{12}(x)}} & \dfrac{5}{12}(x) &\dfrac{7}{12}(x) \\ \hline \\ 5& \boxed{\color{red}{\dfrac{1}{10} . \dfrac{7}{12}(x)}} & \dfrac{57}{120}(x) &\dfrac{63}{120}(x) \\ \hline \\ 6& \boxed{\color{red}{\dfrac{2}{3} . \dfrac{57}{120}(x)}} & \dfrac{95}{120}(x) &\dfrac{25}{120}(x) \\ \hline \\ 7& \boxed{\color{red}{\dfrac{1}{10} . \dfrac{25}{120}(x)}} & \dfrac{975}{1200}(x) &\dfrac{225}{1200}(x) \\ \end{array} \]

ganeshie8 (ganeshie8):

you're correct :) http://prntscr.com/4dd48s

OpenStudy (anonymous):

thank you :)